$SPAD/src/input eq.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool eq.output
)set message test on
)set message auto off
)clear all

———-

Input generated from EquationXmpPage — * —

--S 1 of 12
eq1 := 3*x + 4*y = 5
--R
--R
--R   (1)  4y + 3x= 5
--R                                          Type: Equation(Polynomial(Integer))
--E 1
--S 2 of 12
eq2 := 2*x + 2*y = 3
--R
--R
--R   (2)  2y + 2x= 3
--R                                          Type: Equation(Polynomial(Integer))
--E 2
--S 3 of 12
lhs eq1
--R
--R
--R   (3)  4y + 3x
--R                                                    Type: Polynomial(Integer)
--E 3
--S 4 of 12
rhs eq1
--R
--R
--R   (4)  5

--R                                                    Type: Polynomial(Integer)
--E 4
--S 5 of 12
eq1 + eq2
--R
--R
--R   (5)  6y + 5x= 8
--R                                          Type: Equation(Polynomial(Integer))
--E 5
--S 6 of 12
eq1 * eq2
--R
--R
--R          2             2
--R   (6)  8y  + 14x y + 6x = 15
--R                                          Type: Equation(Polynomial(Integer))
--E 6
--S 7 of 12
2*eq2 - eq1
--R
--R
--R   (7)  x= 1
--R                                          Type: Equation(Polynomial(Integer))
--E 7
--S 8 of 12
eq1**2
--R
--R
--R           2             2
--R   (8)  16y  + 24x y + 9x = 25
--R                                          Type: Equation(Polynomial(Integer))
--E 8
--S 9 of 12
if x+1 = y then "equal" else "unequal"
--R
--R
--R   (9)  "unequal"
--R                                                                 Type: String
--E 9
--S 10 of 12
eqpol := x+1 = y
--R
--R
--R   (10)  x + 1= y

--R                                          Type: Equation(Polynomial(Integer))
--E 10
--S 11 of 12
if eqpol then "equal" else "unequal"
--R
--R
--R   (11)  "unequal"
--R                                                                 Type: String
--E 11
--S 12 of 12
eqpol::Boolean
--R
--R
--R   (12)  false
--R                                                                Type: Boolean
--E 12
)spool
)lisp (bye)

———-

References^{[1] nothing}

6